2. Use Cayley-Hamilton theorem to find A6 − 5A5 + 8A4 − 2A3 − 9A2 + 31A − 36I,

when A=

1 0 3

2 1 −1

1 −1 1


1
Expert's answer
2021-10-11T09:53:43-0400

Find the characteristic polynomial of the matrix A:

"p_A(\lambda)=\det\begin{vmatrix}
1-\lambda & 0 & 3\\
2 & 1-\lambda & -1\\
1 & -1 & 1-\lambda
\end{vmatrix}=(1-\lambda)^3-6-4(1-\lambda)"

"-p_A(\lambda)=\lambda^3-3\lambda^2-\lambda+9"

Cayley-Hamilton theorem claims that "p_A(A)=0", that is, "A^3-3A^2-A+9I=0".

Let's divide "A^6 \u2212 5A^5 + 8A^4 \u2212 2A^3 \u2212 9A^2 + 31A \u2212 36I" by "A^3-3A^2-A+9I" with remainder. We obtain:

"A^6 \u2212 5A^5 + 8A^4 \u2212 2A^3 \u2212 9A^2 + 31A \u2212 36I="

"=(A^3-3A^2-A+9I)(A^3 \u2212 2^2 + 3A \u2212 4I)=0"


Answer. 0.